KNOTS, LINKS, AND 4-MANIFOLDS

RONALD J. STERN RONALD FINTUSHEL

Abstract. We investigate the relationship between isotopy classes of knots and links in S^{3} and the diffeomorphism types of homeomorphic smooth 4-manifolds. As a corollary of this initial investigation, we begin to uncover the surprisingly rich structure of diffeomorphism types of manifolds homeomorphic to the K3 surface.